OpenStudy (anonymous):

Can someone give me an example of a quadratic equation that has an axis of symmetry of x=2 and whose graph opens down.

OpenStudy (anonymous):

y = -x^2 + 4x

OpenStudy (anonymous):

What about the "c" term

OpenStudy (anonymous):

The c term means how far the graph is from the x - axis in terms of the y-axis direction. For example, y = -x^2 + 4x + 3 It means the graph is 3 units upwards of y = -x^2 + 4x and since you just wanted an example, I gave you the simplest, which is y = -x^2 + 4x + 0

OpenStudy (campbell_st):

what about y=(x -1)(x -3) just distribute to get the equation..

OpenStudy (anonymous):

Thanks !

OpenStudy (anonymous):

Here's your medal.